Takealot orders in a certain area can either be delivered by Takealot or picked up by the customer at a pick-up point. Historically, 40% of the orders are delivered by Takealot. Given 8 orders, what is the probability that at least 3 of them will be picked up by the customers?

1
Expert's answer
2021-11-16T14:04:57-0500

for binomial distribution:

"P(x=k)=C^k_np^k(1-p)^{n-k}"


we have:

"p=0.6,n=8"


then:

"P(x\ge3)=1-P(x=0)-P(x=1)-P(x=2)"

"P(x=0)=0.0007"

"P(x=1)=0.0079"

"P(x=2)=0.0413"


"P(x\ge3)=1-0.0007-0.0079-0.0413=0.9501"